# Current Status of Inert Higgs Dark Matter with Dark Fermions

### v1

@@ V Conclusion
We summarize the characteristics of the feasible parameter regions as follows: First, the mass of the scalar DM within the range of $53$ to $74$ GeV remains allowed, regardless of the anomalies in the muon magnetic moment and the $W$ boson mass, as shown in Fig. 5. Second, both the inert Higgs sector and the new Dirac fermions contribute to the $W$ boson mass as illustrated in Fig. 3. Therefore, a smaller charged mass splitting, $\Delta^{\pm}$, is still permissible for explaining the $W$ boson mass excess, which is notable difference from the inert two-Higgs-doublet models Fan:2022dck . Third, the co-annihilation dominant regions with neutral mass splitting $\Delta^{0}\lesssim 10$ GeV are strongly constrained by current and forthcoming LHC searches, as indicated in Fig. 5. Fourth, although the recent LZ results exclude a substantial portion of the parameter space around the Higgs resonance, smaller $\lambda_{S}$ values are still allowed and remain challenging to probe in future DM direct detection experiments. Fifth, there are small regions for DM masses around $63$ GeV and $71-73$ GeV that can simultaneously explain both the GCE and the AMS-02 $\overline{p}$ excess, as seen in the right panel of Fig. 7. Finally, the model parameters $y_{H_{1}}$, $k_{M}$, and $k_{\mu}$, along with the neutral dark fermion mass, are crucial for explaining the muon magnetic moment excess, as shown in the right panel of Fig. 6. However, these parameters are difficult to examine at the LHC. Fortunately, a high-energy muon collider may provide an ideal environment to detect them, potentially confirming or ruling out the predictions in this model.
@@ Acknowledgments
This work was supported by the National Key Research and Development Program of China (No. 2022YFF0503304), and the Project for Young Scientists in Basic Research of the Chinese Academy of Sciences (No. YSBR-092). CTL and XYL are supported by the National Natural Science Foundation of China (NNSFC) under grant No. 12335005 and the Special funds for postdoctoral overseas recruitment, Ministry of Education of China (No. 164080H0262403). This work was also supported in part by the NNSFC under grant Nos. 19Z103010239 and 12350410369, and Vietnam National Foundation for Science and Technology Development (NAFOSTED) under grant number 103.01-2023.50 (VQT). VQT would like to thank the Medium and High Energy Physics group at the Institute of Physics, Academia Sinica for their hospitality during the course of this work.

### v2

@@ V Conclusion
We summarize the characteristics of the feasible parameter regions as follows: First, the mass of the scalar DM within the range of $53$ to $74$ GeV remains allowed, regardless of the anomalies in the muon magnetic moment and the $W$ boson mass, as shown in Fig. 5. Second, both the inert Higgs sector and the new Dirac fermions contribute to the $W$ boson mass as illustrated in Fig. 3. Therefore, a smaller charged mass splitting, $\Delta^{\pm}$, is still permissible for explaining the $W$ boson mass excess, which is notable difference from the inert two-Higgs-doublet models Fan:2022dck . Third, the co-annihilation dominant regions with neutral mass splitting $\Delta^{0}\lesssim 10$ GeV are strongly constrained by current and forthcoming LHC searches, as indicated in Fig. 5. Fourth, although the recent LZ results exclude a substantial portion of the parameter space around the Higgs resonance, smaller $\lambda_{S}$ values are still allowed and remain challenging to probe in future DM direct detection experiments. Fifth, there are small regions for DM masses around $63$ GeV and 717371-73 GeV that can simultaneously explain both the GCE and the AMS-02 $\overline{p}$ excess, as seen in the right panel of Fig. 8. Finally, the model parameters $y_{H_{1}}$, $k_{M}$, and $k_{\mu}$, along with the neutral dark fermion mass, are crucial for explaining the muon magnetic moment excess, as shown in the right panel of Fig. 6. However, these parameters are difficult to examine at the LHC. Fortunately, a high-energy muon collider may provide an ideal environment to detect them, potentially confirming or ruling out the predictions in this model.
@@ Acknowledgments
This work was supported by the National Key Research and Development Program of China (No. 2022YFF0503304), and the Project for Young Scientists in Basic Research of the Chinese Academy of Sciences (No. YSBR-092). CTL and XYL are supported by the National Natural Science Foundation of China (NNSFC) under grant No. 12335005 and the Special funds for postdoctoral overseas recruitment, Ministry of Education of China (No. 164080H0262403). TPT is supported by the Jiangsu Province Post Doctoral Foundation (No. 2024ZB713). This work was also supported in part by the NNSFC under grant Nos. 19Z103010239 and 12350410369, and Vietnam National Foundation for Science and Technology Development (NAFOSTED) under grant number 103.01-2023.50 (VQT). VQT would like to thank the Medium and High Energy Physics group at the Institute of Physics, Academia Sinica for their hospitality during the course of this work.